okay heres my problem ive tried to solve it but it keeps coming out wrong....
18x - (8x - 7) = 67

and also this one....
11 ( 4 - 6y) + 5 ( 13y + 11) = 9

Answer by doukungfoo(195)

Ok first use the Distributive Property to get rid of the parentheses so
18x - (8x - 7) = 67
multiply -1 times 8x and -1 times -7 to get
18x - 8x + 7 = 67
Now combine like terms
10x + 7 = 67
Next subtract 7 from both sides of the equal sign
10x + 7 - 7 = 67 - 7
10x = 60
Finally to solve for x divide each side of the equal sign by 10
10x/10 = 60/10
x = 6
To check just plug 6 into the original equation
-----------------------------------------------------------------------------
The second problem also starts the Distributive Property
11 ( 4 - 6y) + 5 ( 13y + 11) = 9
First 11 is multiplied by 4 and -6y
44 - 66y + 5 ( 13y + 11) = 9
Then 5 is multiplied by 13y and 11
44 - 66y + 65y + 55 = 9
Now combine like terms
99 - 1y = 9
add 1y to both sides
99 -1y + 1y = 9 + 1y
99 = 9 + y
subtract 9 from both sides
99 - 9 = 9 - 9 + y
90 = y